OpenStudy (anonymous):

Let s = 12. What is the value of 27 + 3s?

OpenStudy (anonymous):

@flower2468

OpenStudy (jjcoolja12):

it may not be right but this is what i got (3)(12)+27 =36+27 =63

OpenStudy (xx_gamergirl_xx):

@jjcoolja12 please don't give out direct awnsers.

OpenStudy (mathmale):

You are asked to evaluate 27 + 3s when x=12. Bet you've done this before. show you work.

OpenStudy (mathmale):

No direct answers, please. If you can explain the steps, explain the steps and let the other person find his or her own answers.

OpenStudy (jjcoolja12):

ok i think i get what you mean now